OpenStudy (anonymous):

Alex graphed the system of equations and stated that the solution is (5, 2). Check his solution algebraically. Which statement is true? A. The graph for is incorrect. B. The graph for is incorrect. C. (5, 2) is the solution to the system of equations. D. (2, 5) is the solution to the system of equations.

OpenStudy (anonymous):

@Nnesha @undeadknight26 @HaileyD

OpenStudy (anonymous):

awnser is D

Nnesha (nnesha):

you have two equation this question is same like last one

Nnesha (nnesha):

and another way is to ( 5 , 2) use this point and plug that into equation if both side are equal then yes this is your answer if not then incorrect

OpenStudy (anonymous):

@CrystalBurley how did you find the answer

OpenStudy (anonymous):

@Nnesha that was wrong

OpenStudy (anonymous):

ok how I got awnser is here The solution is where the lines cross...so the solution is (5,2)...if the lines are graphed right lets check your lines y = -3/5x + 5 --- y intercept is (0,5) to find x intercept sub in 0 for y 0 = -3/5x + 5 3/5x = 5 x = 5/(3/5) x = 5*5/3 x = 25/3 = 8 1/3....x intercept is (25/3,0) y = 2/5x + 0 --- y int is (0,0) 0 = 2/5x + 0 -2/5x = 0 x int = 0 Its not A, B, or C...so I am going with D

Nnesha (nnesha):

@bazel_e what's wrong ????

Nnesha (nnesha):

i didn't give you the answer so hows something wrong ?????@ bazel_e

OpenStudy (anonymous):

L-O-L